ABX is estimating its WACC. The company has collected the following information:

▪ Its capital structure consists of 40 percent debt and 60 percent common equity.

▪ The company has 20-year bonds outstanding with a 9 percent annual coupon that are trading at

 par.

▪ The company’s tax rate is 40 percent.

▪ The risk-free rate is 5.5 percent.

▪ The market risk premium is 5 percent.

▪ The stock’s beta is 1.4.

Calculate the company’s WACC.


Anyone can help me answer
Answer #1

After tax cost of debt(kd) = 9%(1 - 0.40) = 5.40%

According to the CAPM,

Ke = Rf + beta(Market Risk Premium)

= 5.5% + 1.4(5%) = 12.50%

WACC = Wd*Kd + We*Ke

= (0.4*5.40%) + (0.6*12.50%)

= 2.16% + 7.50% = 9.66%
Answer #2

Using formulas in Chap9 of Funda of Fina Mgt by Brigham&Houston, we get
WACC=wd*kd(1-T) + wp*kp +wc*ks. Here there is no preferd stock. Sokp=0
wd=40% debt = 0.4
wc= 60% equity = 0.6

ks=krf+(km-krf)beta = 5.5 + 5*1.4=5.5+7=12.7%
Cost of Debt = kd(1-T) = 9(1-40%)= 9*0.6=5.4%

So WACC = 0.4*5.4 + 0.6*12.7 = 2.16+7.62 = 9.78%